the pendulum bob is 0.1m from the floor. if the pendulum is released from point A, what will be the speed of the bob at point B? How high will the pendulum bob (from the floor) be at point C if the total mechanical energy is conserved?
1
Expert's answer
2020-03-09T10:57:54-0400

Consider the following situation:



If the point A is 0.1 m higher than the point B (the lowest point of the trajectory), the speed can be expressed from energy conservation law:

"mgh=\frac{1}{2}mv^2,\\
\space\\v=\sqrt{2gh}=\sqrt{2\cdot9.8\cdot0.1}=1.4\text{ m/s}."

We don't even need to know the mass!

Since the total mechanical energy is conserved, the bob will reach the same height at point C after passing the bottom B.